A constant volume of gas thermometer is used to measure the room temperature during an experiment.If the pressure at 273k is 220mmHg and 290mmHg at 373k.What is room temperature if the thermometer records pressure of 234mmHg?
1
Expert's answer
2021-10-15T10:23:20-0400

Determine the rate of change in pressure with temperature:


"r=\frac{P_2-P_1}{T_2-T_1}=\frac{290-220}{373-273}=0.7\text{ mmHg/K}."

At 234 mmHg:


"r=\frac{P_3-P_1}{T_3-T_1},\\\space\\
T_3=\frac{P_3-P_1}{r}+T_1=293\text{ K}."